Problem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure [LO5-1, LO5-3, LO5-4, LO5-5, LO5-6]

Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,300 units × $20 per unit) $ 266,000
Variable expenses 159,600
Contribution margin 106,400
Fixed expenses 118,400
Net operating loss $ (12,000 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,300 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $80,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $39,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.80 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $5,000?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $53,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,300 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)?

Compute the company’s CM ratio and its break-even point in unit sales and dollar sales. (Do not round intermediate calculations. Round "CM ratio" to the nearest whole percentage (i.e., 0.234 should be entered as "23").)

CM ratio %
Break-even point in unit sales
Break-even point in dollar sales

The president believes that a $6,300 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $80,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income? (Do not round intermediate calculations.)

by

Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $39,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? (Losses should be entered as a negative value.)

Revised net operating income (loss)

Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.80 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $5,000? (Do not round intermediate calculations. Round final answer to the nearest whole unit.)

Show less

Unit sales to attain target profit

Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $53,000 each month. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. (Do not round intermediate calculations. Round "CM ratio" to the nearest whole percentage (i.e., 0.234 should be entered as "23") and other answers to the nearest whole number.)

Show less

CM ratio %
Break-even point in unit sales
Break-even point in dollar sales

Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $53,000 each month. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. (Do not round intermediate calculations. Round "CM ratio" to the nearest whole percentage (i.e., 0.234 should be entered as "23") and other answers to the nearest whole number.)

Show less

CM ratio %
Break-even point in unit sales
Break-even point in dollar sales

Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $53,000 each month. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)?

yes

or

no
Answer #1
contribution margin per unit= 106400/13300
8
1) CM ratio = contribution/sales
106400/266000
40.00%
BEP(units) = total fixed cost/contribution margin per unit
118400/8
14800
BEP(dollars) = 14800*20
296000
CM ratio 40%
Break even point in units 14800
Break even point in dollars 296000
2) increase in contribution (80000*40%) 32000
less : increase in advertising budget 6,300
increase in net income 25,700
increases by 25,700
3) units = 13300*2 = 26600 units ; selling price = 20*90%=$18
Contribution Income statement
Sales (26600*18) 478800
Variable expense (26600*12) 319200
Contribution margin 159600
Fixed expenses (118400+39000) 157,400
Net income 2,200
4) New contribution margin = 8-.80
7.2
BEP(units) = (total fixed cost+target profit)/contribution per unit
(118400+5000)/7.2
17138.89
Sales units 17,139
5)
CM ratio = contribution/sales
11./20
55.00%
BEP(units) = total fixed cost/contribution margin per unit
171400/11
15582
BEP(dollars) = 171400/55%
311636
CM ratio 55%
Break even point in units 15582
Break even point in dollars 311636
20300
b) Not Automated Automated
total per unit % total per unit %
Sales 406000 20 100% 406000 20 100%
Variable expenses 243600 12 60% 182700 9 45%
Contribution margin 166400 8 40% 223300 11 55%
Fixed expenses 118,400 171,400
Net operating income 48,000 51,900